### v1

# GINO-Q: Learning an Asymptotically Optimal Index Policy for Restless Multi-armed Bandits

Abstract
The restless multi-armed bandit (RMAB) framework is a popular model with applications across a wide variety of fields. However, its solution is hindered by the exponentially growing state space (with respect to the number of arms) and the combinatorial action space, making traditional reinforcement learning methods infeasible for large-scale instances. In this paper, we propose GINO-Q, a three-timescale stochastic approximation algorithm designed to learn an asymptotically optimal index policy for RMABs. GINO-Q mitigates the curse of dimensionality by decomposing the RMAB into a series of subproblems, each with the same dimension as a single arm, ensuring that complexity increases linearly with the number of arms. Unlike recently developed Whittle-index-based algorithms, GINO-Q does not require RMABs to be indexable, enhancing its flexibility and applicability. Our experimental results demonstrate that GINO-Q consistently learns near-optimal policies, even for non-indexable RMABs where Whittle-index-based algorithms perform poorly, and it converges significantly faster than existing baselines.
Introduction
A restless multi-armed bandit (RMAB) models a sequential decision-making problem, in which a set of resources must be allocated to $N$ out of $M$ ($1\leq N<M$) arms at each discrete time step. Here, each arm represents a dynamic process. Upon the resource allocation, each arm generates a reward and may transition to a new state. These arms evolve independently except for simultaneously being subject to the resource constraint. Hence, RMABs essentially constitute a special case of weakly coupled dynamic programs (Adelman and Mersereau 2008). The objective is to determine an optimal policy for selecting the arms at each time step in a way that maximizes the expected reward over the entire time horizon. RMABs find applications in diverse fields, including resource allocation (Wang et al. 2019), opportunistic scheduling (Wang and Chen 2021), public health interventions (Mate et al. 2022), and many more.
This paper investigates the design of an efficient reinforcement learning (RL) algorithm for RMABs. The curse of dimensionality is a key challenge in solving dynamic programs. Unfortunately, this problem is particularly severe in RMABs due to the exponential growth of RMAB’s dimension with the number of arms. In fact, Papadimitriou and Tsitsiklis (1999) proved that RMABs are P-SPACE hard. As a result, directly treating RMABs as Markov decision processes (MDPs) and applying RL algorithms is inefficient, and, in some cases, infeasible—particularly for large-scale RMABs. For instance, consider a moderate-scale RMAB with $M=100$ and $N=25$. Suppose each arm exhibits 10 states. Then the RMAB encompasses an astronomical $10^{100}$ possible states, along with ($100\choose 25$ $\approx 2.4\times 10^{23}$ valid actions. Such complexity presents a great challenge for conventional RL algorithms.
To address this challenge, we propose an efficient algorithm that decomposes the RMAB into a series of subproblems and solve them using a three-timescale stochastic approximation algorithm. Our method draws inspiration from a recently developed planning algorithm for RMABs known as the gain index policy (Chen and Liew 2024). In a nutshell, the gain index policy computes a gain index for each arm at each time step based on the arm’s current state. Then it selects the top $N$ arms with the highest gain indices. The gain index, defined in terms of the Q function of an MDP associated with each arm, quantifies the advantage of selecting an arm in its current state compared to not selecting it. Our algorithm, referred to as Gain-Index-Oriented Q (GINO-Q) learning, aims to efficiently learn this gain index.
Related Work    The success of deep RL has attracted considerable research interest in its application to practical problems that can be modeled as RMABs. Notable examples include wireless sensor scheduling (Leong et al. 2020), dynamic multichannel access (Wang et al. 2018; Demirel et al. 2018), intelligent building management (Wei, Wang, and Zhu 2017). However, a common limitation across these studies is the poor scalability of deep RL in RMABs. The experimental results presented in these papers are limited to small-scale scenarios, with the number of arms $M<10$.
In the planning scenario, where the system knowledge is known, the Whittle index policy (Whittle 1988) is recognized as one of the most efficient heuristic algorithms for addressing RMAB problems. The Whittle index policy computes an index for each state of each arm based solely on the arm’s parameters. This has inspired a series of studies focused on applying RL algorithms to learn Whittle indices. Fu et al. (2019) proposed a Q-learning algorithm to learn the Whittle indices; however, their experiments revealed that the proposed algorithm struggles to accurately learn the Whittle indices. Subsequently, Avrachenkov and Borkar (2022) introduced Whittle-Index-Based Q-learning (WIBQ), a two-timescale stochastic approximation algorithm designed to learn the Whittle indices. They proved theoretically that WIBQ converges to the Whittle indices for indexable RMABs. Further advancing this work, Xiong and Li (2023) enhanced WIBQ to handle arms with large state spaces by coupling WIBQ with neural network function approximation. Additionally, Nakhleh et al. (2021) converted the computation of Whittle indices to an optimal control problem and proposed Neurwin, a neural network-based approach for computing Whittle indices.
In practical applications, Whittle-index-based learning algorithms have been employed in adaptive video streaming (Xiong et al. 2022), wireless edge caching (Xiong et al. 2024), and preventive healthcare (Biswas et al. 2021). Moreover, the Whittle index policy has also inspired studies on online learning for RMABs (Wang et al. 2023; Xiong, Wang, and Li 2022; Xiong, Li, and Singh 2022).
A limitation of the Whittle index policy is that it applies only to the RMABs that possess the “indexability” property—a condition that does not naturally hold for all RMABs. While some studies have explored sufficient conditions for the indexability of RMABs (Niño-Mora 2001, 2007; Glazebrook, Ruiz-Hernandez, and Kirkbride 2006), these conditions are stringent and are satisfied only by a limited subset of RMABs. In most cases, determining the indexability of an RMAB is a difficult task that requires a rigorous analytical effort reliant on the knowledge of the system parameters (Chen, Liew, and Shao 2022; Liu and Zhao 2010; Villar 2016). Consequently, verifying the indexability is usually unachievable without full knowledge of the system. As will be shown in the experiments section, applying the Whittle index policy to a non-indexable RMAB can result in arbitrarily poor performance outcomes. Therefore, the use of Whittle-index-based learning algorithms poses a significant risk when applied to RMABs that lack indexability guarantees. In contrast, the gain index policy eliminates the need for indexability and can be applied to any RMAB.
Main Contributions    In this paper, we introduce GINO-Q learning, a three-timescale stochastic approximation algorithm designed to learn gain indices for RMABs. The gain index policy is asymptotically optimal in the number of arms; however, computing gain indices is a computationally intensive task even when the system parameters are known (Chen and Liew 2024). Our novel approach integrates Q-learning, SARSA learning, and stochastic gradient descent across different timescales to efficiently learn gain indices in the model-free setting. The key advantages of GINO-Q are: (1) Scalability: By decomposing the RMAB into a series of subproblems, GINO-Q only needs to solve problems with the same dimension as a single arm, avoiding astronomical state and action spaces and ensuring that the complexity increases linearly with $M$ rather than exponentially. (2) Flexibility: GINO-Q does not require the RMAB to be indexable, broadening its applicability. Our experiments demonstrate that Whittle-index-based learning algorithms can perform poorly in non-indexable RMABs. In contrast, GINO-Q consistently learns near-optimal policies, even for non-indexable RMABs.
Notation    For any positive integer $M$, we will use $[M]$ to denote the set of positive integers between 1 and $M$, i.e., $[M]=\{1,2,\cdots,M\}$. $\mathbb{E}[\cdot]$ denotes the expectation.
Problem Statement and Preliminaries
An RMAB consists of $M$ arms $\{\mathcal{B}_{i}:i\in[M]\}$. Each arm $\mathcal{B}_{i}$ is an MDP represented by a tuple $(\mathcal{S}_{i},\mathcal{A}_{i},r_{i},p_{i})$, where $\mathcal{S}_{i}$ is the state space, $\mathcal{A}_{i}$ is the action space, $r_{i}:\mathcal{S}_{i}\times\mathcal{A}_{i}\to\mathbb{R}$ is the reward function, and $p_{i}$ is the transition kernel. Let $s_{i}^{t}$ and $a^{t}_{i}$ denote the state and action of $\mathcal{B}_{i}$ at time $t$. In particular, we have $\mathcal{A}_{i}=\{0,1\}$ for all $i\in[M]$. We will say that arm $\mathcal{B}_{i}$ is activated at time $t$ if $a^{t}_{i}=1$. At any time step, all arms evolve independently based on their actions; that is,
$P(s^{t+1}_{1},s^{t+1}_{2},\cdots,s^{t+1}_{M}|s^{t}_{1},s^{t}_{2},\cdots,s^{t}_{M},a^{t}_{1},a^{t}_{2},\cdots,a^{t}_{M})$
$=$ $\prod_{i=1}^{M}p_{i}(s^{t+1}_{i}|s^{t}_{i},a^{t}_{i}).$
However, there is a constraint on the actions that weakly couples all the arms. Specifically, at each time step, only $N$ arms can be activated ($1\leq N<M$). That is, $\sum_{i=1}^{M}a^{t}_{i}=N$ for all $t$. The objective is to identify an optimal policy that maximizes the cumulative reward obtained from all the arms. Mathematically, an RMAB is a constrained optimization problem defined as follows:
$\max_{\{a^{t}_{i}:i\in[M],t\geq 1\}}$ $\lim\limits_{T\to\infty}\frac{1}{T}\mathbb{E}\left[\sum_{t=1}^{T}\sum_{i=1}^{M}r_{i}\left(s^{t}_{i},a^{t}_{i}\right)\right]$ (1)
subject to $\sum_{i=1}^{M}a^{t}_{i}=N,\quad\forall t.$ (2)
Throughout this paper, we assume that all arms are unichain MDPs. As a result, the objective function (1) is always well-defined and independent of the initial state.
The Gain Index Policy
An RMAB can be formulated as an MDP with state $(s^{t}_{1},s^{t}_{2},\cdots,s^{t}_{M})$ and action $(a^{t}_{1},a^{t}_{2},\cdots,a^{t}_{M})$. However, the state and action spaces of this MDP grow exponentially with $M$. As a result, the problem becomes challenging to solve when $M$ is large. Since the $M$ arms are weakly coupled only via constraint (2), a common approach to mitigate the complexity is to decompose the RMAB into single-arm problems by relaxing the constraint. Building on this concept, Chen and Liew (2024) introduced a gain index policy and proved its asymptotic optimality.
In particular, we can relax constraint (2) of the RMAB problem to the following:
$\lim\limits_{T\to\infty}\frac{1}{T}\mathbb{E}\left[\sum_{t=1}^{T}\sum_{i=1}^{M}a^{t}_{i}\right]=N.$ (3)
Replacing (2) by (3) leads to a relaxed RMAB. We can further transform the relaxed problem using the Lagrange multiplier method and yield an unconstrained problem:
$\max_{\{a^{t}_{i}\}}\inf_{\lambda}\ \lim\limits_{T\to\infty}\frac{1}{T}\mathbb{E}\left[\sum_{t=1}^{T}\sum_{i=1}^{M}\left[r_{i}\left(s^{t}_{i},a^{t}_{i}\right)-\lambda a^{t}_{i}\right]\right]+{N\lambda},$ (4)
where $\lambda\in\mathbb{R}$ is the Lagrange multiplier. The max and inf in (4) can be interchanged. That is, (4) is equivalent to:
$\inf_{\lambda}\max_{\{a^{t}_{i}\}}\ \lim\limits_{T\to\infty}\frac{1}{T}\mathbb{E}\left[\sum_{t=1}^{T}\sum_{i=1}^{M}\left[r_{i}\left(s^{t}_{i},a^{t}_{i}\right)-\lambda a^{t}_{i}\right]\right]+{N\lambda}.$ (5)
For any fixed $\lambda$, (5) can be decoupled into $M$ subproblems:
$J_{i}(\lambda):\max_{\{a^{t}_{i}:t\geq 1\}}\ \lim\limits_{T\to\infty}\frac{1}{T}\mathbb{E}\left[\sum_{t=1}^{T}\left[r_{i}\left(s^{t}_{i},a^{t}_{i}\right)-\lambda a^{t}_{i}\right]\right],$ (6)
where $i\in[M]$. We refer to each $J_{i}(\lambda)$ as a single-arm problem. Note that $J_{i}(\lambda)$ is an MDP associated with $\mathcal{B}_{i}$: they share the same state space $\mathcal{S}_{i}$, action space $\mathcal{A}_{i}$, and transition kernel $p_{i}$. However, there is a distinction between them in terms of their reward functions. The reward function of $J_{i}(\lambda)$ is defined as $r_{i}(s,a)-\lambda a$, where $\lambda$ can be viewed as the cost of action $a=1$. If we denote the optimal value of problem $J_{i}(\lambda)$ by $g_{i}(\lambda)$, it can be determined by the Bellman equation as follows (Puterman 2014):
$V_{i}(s,\lambda)+g_{i}(\lambda)=\max_{a\in\{0,1\}}\left\{Q_{i}(s,a,\lambda)\right\},\ s\in\mathcal{S}_{i},$ (7)
where $V_{i}(s,\lambda)$ is the state value function and $Q_{i}(s,a,\lambda)$ is the state-action value function:
$Q_{i}(s,a,\lambda)\triangleq r_{i}(s,a)-\lambda a+\sum_{s^{\prime}\in\mathcal{S}_{i}}p_{i}(s^{\prime}|s,a)V_{i}(s^{\prime},\lambda).$
Here, we express $V_{i}$ and $Q_{i}$ as functions of $\lambda$ to highlight their dependence on $\lambda$. Then, (5) reduces to
$\inf_{\lambda}\ f(\lambda)\triangleq\sum_{i=1}^{M}g_{i}(\lambda)+N\lambda.$ (8)
Denoting by $\lambda^{*}$ the optimal solution to problem (8), then gain index policy is defined as follows:
Definition 1 (Gain index policy).
For each arm $i\in[M]$ and each state $s\in\mathcal{S}_{i}$, a gain index is defined as:
$W_{i}(s)\triangleq Q_{i}(s,1,\lambda^{*})-Q_{i}(s,0,\lambda^{*}).$ (9)
Then the gain index of the $i$-th arm at time $t$ is given by $W_{i}(s^{t}_{i})$. The gain index policy activates the $N$ arms with the largest $N$ gain indices, with ties broken arbitrarily.
Intuitively, the gain index $W_{i}(s)$ evaluates the “gain” of activating arm $\mathcal{B}_{i}$ when it is in state $s$. It turns out that this heuristic algorithm is asymptotically optimal in $M$ under some mild conditions. Two arms $\mathcal{B}_{i}$ and $\mathcal{B}_{j}$ are considered to belong to the same class if they share the same MDP model (i.e., having the same state space, reward function, and transition kernel). In this paper, we assume that the $M$ arms can be grouped into $K$ classes, with the proportion of the arms in the $k$-th class denoted by $\eta_{k}\in(0,1]$, where $\sum_{k=1}^{K}\eta_{k}=1$. The following result is proved in (Chen and Liew 2024):
Theorem 1.
If $K<\infty$ is fixed, and $\{\eta_{k}\}$ and $N/M$ are fixed. Then the gain index policy is asymptotically optimal in the following sense:
$\lim_{M\to\infty}\frac{1}{M}G_{M}^{ind}=\lim_{M\to\infty}\frac{1}{M}G_{M}^{opt},$
where $G_{M}^{ind}$ and $G_{M}^{opt}$ denote the cumulative reward of the RMAB under the gain index policy and the optimal policy, respectively.
Whittle Index Policy and Indexability
In addition to the gain index policy, the renowned Whittle index policy (Whittle 1988) is also an effective approach for RMABs. It shares the same asymptotic optimality as described in Theorem 1 (Weber and Weiss 1990). Similar to the gain index policy, the Whittle index policy assigns an index to each state of each arm and selects at each time the top $N$ arms with the highest indices for activation. Whittle’s index for state $s$ of arm $\mathcal{B}_{i}$ is defined as the value of $\lambda$ that ensures equal optimality between taking action 0 and action 1 in state $s$ within the single-arm problem $J_{i}(\lambda)$.
A limitation of the Whittle index policy is its applicability only to indexable RMABs. Specifically, we denote the set of states in which action 0 is optimal for problem $J_{i}(\lambda)$ by $\mathcal{E}_{i}(\lambda)$. An arm $\mathcal{B}_{i}$ is considered indexable if $\mathcal{E}_{i}(\lambda)$ expands monotonically from the empty set to the entire state space $\mathcal{S}_{i}$ as $\lambda$ increases from $\infty$- ∞ to $\infty$. An RMAB is indexable if all its arms are indexable. The indexability of an arm ensures the uniqueness of the Whittle index for each state. In the case of non-indexable RMABs, there can be multiple values of $\lambda$ where both actions are equally optimal in a state, leading to confusions in defining the Whittle index.
Despite sharing the same asymptotic optimality, the gain index policy does not require indexability, making it applicable to all RMABs. This advantage is a key reason we advocate for learning gain indices instead of Whittle indices.
Gain-Index-Oriented Q Learning
This section presents the GINO-Q learning algorithm for RMABs. Our approach involves decomposing the RMAB into single-arm problems and learning the gain indices for each arm. The main challenge in learning these gain indices lies in determining the optimal solution to (8), i.e., $\lambda^{*}$. Once $\lambda^{*}$ is determined, the gain indices for any arm $\mathcal{B}_{i}$ can be derived from the Q function of problem $J_{i}(\lambda^{*})$, which can be learned using conventional Q learning algorithms.
To address this challenge, we propose a three-timescale stochastic approximation algorithm. This algorithm utilizes different timescales to update $\lambda$ and the Q function of $J_{i}(\lambda)$ at slow and medium rates, respectively. Additionally, the fast timescale is designed to estimate the derivative of $g_{i}(\lambda)$ with respect to $\lambda$, which is crucial for updating $\lambda$.
We will begin by introducing some useful definitions and establishing key properties of the optimization problem (8), which form the basis of our method. Let $\pi$ be a policy for the single-arm problem $J_{i}(\lambda)$, and $V_{i}^{\pi}(s)$ and $g_{i}^{\pi}$ be the associated value function and long-term average reward, respectively. Define
$h_{i}^{\pi}(s)\triangleq\lim\limits_{T\to\infty}\frac{1}{T}\mathbb{E}_{\pi}\left[\sum_{t=1}^{T}a_{i}^{t}|s^{1}_{i}=s\right],\ \forall s\in\mathcal{S}_{i}.$
Since all the arms are unichain, $h_{i}^{\pi}(s)$ is invariant with $s$ for all $i$. Hence, we will simply write $h_{i}^{\pi}$ in the following. It is easy to see that ${dg_{i}^{\pi}}/{d\lambda}=-h_{i}^{\pi}$. In fact, let $c(s,a)=a$ be the cost of taking action $a$ in state $s$, $\forall s\in\mathcal{S}_{i},a\in\{0,1\}$. Then $h_{i}^{\pi}$ can be viewed as the long-term average cost of MDP $(\mathcal{S}_{i},\mathcal{A}_{i},c,p_{i})$ under policy $\pi$. It can be determined by the following Bellman equation:
$D_{i}^{\pi}(s,a)+h_{i}^{\pi}=a+\sum_{s^{\prime}\in\mathcal{S}_{i}}\sum_{a^{\prime}\in\mathcal{A}}p_{i}(s^{\prime}|s,a)\pi(a^{\prime}|s^{\prime})D_{i}^{\pi}(s^{\prime},a^{\prime}),$
where $D^{\pi}_{i}(s,a)$ is the state-action value function of MDP $(\mathcal{S}_{i},\mathcal{A},c,p_{i})$ under policy $\pi$, and $\pi(a^{\prime}|s^{\prime})$ is the probability of taking actin $a^{\prime}$ in state $s^{\prime}$.
Let $\pi_{i}^{\lambda}$ denote the optimal policy for problem $J_{i}(\lambda)$, then
$f^{\prime}(\lambda)=\sum_{i=1}^{M}\frac{dg_{i}(\lambda)}{d\lambda}+N=N-\sum_{i=1}^{M}h^{\pi_{i}^{\lambda}}_{i},$ (10)
where $f^{\prime}(\lambda)$ is the derivative of function $f(\lambda)$. Drawing from these concepts, we can establish the following properties regarding $f(\lambda)$ and $\lambda^{*}$:
Lemma 1.
For any RMAB with bounded reward functions $\{r_{i}\}_{i\in[M]}$, $f(\lambda)$ is a piecewise linear and convex function. In addition, there always exist a bounded $\lambda^{*}$ that achieves the minimum value of $f(\lambda)$.
Proof.
By definition, we have $g_{i}(\lambda)=\max_{\pi}g^{\pi}_{i}(\lambda)$. As mentioned above, for any policy $\pi$, the derivative ${dg_{i}^{\pi}}/{d\lambda}=-h_{i}^{\pi}\leq 0$ is independent of $\lambda$, implying $g^{\pi}_{i}(\lambda)$ is a decreasing linear function of $\lambda$. Consequently, $g_{i}(\lambda)$ is decreasing, piecewise linear and convex given that it is the maximum of a set of decreasing linear functions. It then follows immediately that $f(\lambda)$ is piecewise linear and convex.
Furthermore, since $a^{t}_{i}\in\{0,1\}$ for all $t$ and $i$, we have $0\leq h^{\pi}_{i}\leq 1$. Hence $-1\leq{dg_{i}}/{d\lambda}\leq 0$ for all $i\in[M]$. It is easy to verify that, as $\lambda\to-\infty$, the optimal policy for $J_{i}(\lambda)$ is taking action 1 all the time. That is,
$\lim\limits_{\lambda\to-\infty}\frac{dg_{i}(\lambda)}{d\lambda}=-1,\ \forall i\in[M].$
Using the same argument yields $\lim_{\lambda\to\infty}{dg_{i}}/{d\lambda}=0$. It follows that
$\lim\limits_{\lambda\to-\infty}f^{\prime}(\lambda)=N-M<0,\ \lim\limits_{\lambda\to\infty}f^{\prime}(\lambda)=N>0.$
Therefore, there must exist a bounded $\lambda_{0}$ satisfying
$f^{\prime}(\lambda^{-}_{0})\leq 0,\ f^{\prime}(\lambda^{+}_{0})\geq 0.$
This $\lambda_{0}=\lambda^{*}$ achieves the minimum of $f(\lambda)$. ∎
Our goal is to learn $\lambda^{*}$ and the associated Q function of problem $J_{i}(\lambda^{*})$ for all $i\in[M]$. To achieve this, we propose a simple but effective algorithm for updating $\lambda$ and learning the associated Q function. In particular, we fix a stepsize sequence $\{\theta^{t}:t\geq 1\}$, and employ the stochastic gradient-descent method to update $\lambda$:
$\lambda^{t+1}=\lambda^{t}-\theta^{t}\hat{f}^{\prime}(\lambda^{t}),$ (11)
where $\hat{f}^{\prime}(\lambda^{t})$ is an estimator of $f^{\prime}(\lambda^{t})$. Meanwhile, we use the standard relative value iteration (RVI) Q learning algorithm (Abounadi, Bertsekas, and Borkar 2001) to learn the Q function of every single-arm problem $J_{i}(\lambda^{t})$ for the current $\lambda^{t}$. That is, $\forall i\in[M]$:
$Q_{i}^{t+1}(s^{t}_{i},a^{t}_{i})=$ $Q_{i}^{t}(s^{t}_{i},a^{t}_{i})+\beta^{t}_{i}[r_{i}(s^{t}_{i},a^{t}_{i})-\lambda^{t}a^{t}_{i}+$
$\max_{a}Q_{i}^{t}(s^{t+1}_{i},a)-Q_{i}^{t}(s^{t}_{i},a^{t}_{i})-g^{t}_{i}],$ (12)
where $\{\beta^{t}_{i}:t\geq 1\}$ is a predefined stepsize sequence. While $g_{i}^{t}$ can be estimated in various ways (Abounadi, Bertsekas, and Borkar 2001), one of the most widely used methods is:
$g_{i}^{t}=\frac{1}{2|\mathcal{S}|}\sum_{s\in\mathcal{S}_{i}}\sum_{a\in\mathcal{A}_{i}}Q^{t}_{i}(s,a).$
We proceed by constructing the estimator $\hat{f}^{\prime}(\lambda)$ according to (10). Given any $\lambda$, the optimal policy $\pi^{\lambda}_{i}$ for $J_{i}(\lambda)$ selects actions greedily according to the optimal Q function of $J_{i}(\lambda)$, denoted by $Q_{i}$. At time step $t$, $Q^{t}_{i}$ serves as an estimate of $Q_{i}$. Hence, the policy that select actions greedily according to $Q^{t}_{i}$ can be regarded as an estimate of $\pi^{\lambda}_{i}$. To learn the average cost of this policy for MDP $(\mathcal{S}_{i},\mathcal{A}_{i},c,p_{i})$, we employ the SARSA algorithm (Sutton and Barto 2018):
$D_{i}^{t+1}(s^{t}_{i},a^{t}_{i})=D_{i}^{t}(s^{t}_{i},a^{t}_{i})+\alpha^{t}_{i}[$ $a^{t}_{i}+D_{i}^{t}(s^{t+1}_{i},a_{i}^{t+1})-$
$D_{i}^{t}(s^{t}_{i},a^{t}_{i})-h^{t}_{i}],$ (13)
where $\{\alpha^{t}_{i}:t\geq 1\}$ is the stepsize sequence, action $a_{i}^{t+1}$ is selected greedily based on $Q^{t}_{i}$, that is, $a_{i}^{t+1}=\arg\max_{a}Q_{i}^{t}(s^{t+1}_{i},a)$. The average cost is computed by
$h_{i}^{t}=\frac{1}{2|\mathcal{S}|}\sum_{s\in\mathcal{S}_{i}}\sum_{a\in\mathcal{A}_{i}}D^{t}_{i}(s,a).$
Then $f^{\prime}(\lambda^{t})$ can be estimated by $\hat{f}^{\prime}(\lambda^{t})=N-\sum_{i=1}^{M}h^{t}_{i}$.
We define the stepsize sequences of the three coupled iterates properly so that they form a three-timescale stochastic approximation algorithm, with updates (11), (Gain-Index-Oriented Q Learning), and (Gain-Index-Oriented Q Learning) operating in the slow, medium, and fast timescales, respectively. In particular, define [^1] In practice, $t$ in the expressions of $\alpha^{t}_{i}$ and $\beta^{t}_{i}$ can be replaced by $n(s^{t}_{i},a^{t}_{i})$, the number of occurrences of $(s^{t}_{i},a^{t}_{i})$ up to time $t$. If $\lambda^{t}$ changes greatly compared with $\lambda^{t-1}$, we may reset $n(s,a)$ as if encountering a new RL problem. This may speed up convergence.
$\alpha^{t}_{i}=\frac{C_{1}}{t},\beta^{t}_{i}=\frac{C_{2}}{t\sqrt{\log t}},\theta^{t}=\frac{C_{3}}{t\log t}\mathbf{1}\{t(\text{mod }C_{4})=0\},$
where $\{C_{k}\}$ are constants, $\mathbf{1}\{\cdot\}$ is the indicator function, $\{\alpha^{t}_{i}\}$ and $\{\beta^{t}_{i}\}$ are invariant across $i\in[M]$. We can verify that $\sum_{t=1}^{\infty}x^{t}=\infty$ and $\sum_{t=1}^{\infty}(x^{t})^{2}<\infty$ for $x=\alpha_{i},\beta_{i}$ or $\theta$. Update (Gain-Index-Oriented Q Learning) is faster than (Gain-Index-Oriented Q Learning) because $\beta^{t}_{i}=o(\alpha_{i}^{t})$. A similar argument indicates that update (Gain-Index-Oriented Q Learning) is faster than (11). The GINO-Q algorithm is summarized in Algorithm 1. A detail to note is line 12, where $\lambda^{t}$ is updated only if the estimated derivative is decreased in absolute value. This trick is justified by the convexity of $f(\lambda)$.
Remark: If the learner has knowledge of the arm classification (but not the specific parameters of arms), updates (Gain-Index-Oriented Q Learning) and (Gain-Index-Oriented Q Learning) do not need to be performed for each individual arm. Instead, it is sufficient to maintain a pair of $(Q^{t}_{i},D^{t}_{i})$ for each class. This is because arms within the same class share the same Q and D functions. As a result, the complexity of GINO-Q increases linearly with respect to the number of classes $K$. If the arm classification is unknown, then the complexity scales linearly with respect to the number of arms $M$, which still represents a significant reduction compared to the exponential growth of the state space.
Algorithm 1 GINO-Q Learning
Input: Integer $T$, learning rates $\{\alpha^{t}_{i}\}$, $\{\beta^{t}_{i}\}$, and $\{\theta^{t}\}$
Initialization: Let $t=1,\lambda^{1}=0,y^{0}=M$. Reset the RMAB and receive the initial arm states $\{s^{1}_{i}\}$
Output: Gain indices of all arms
1:  while $t\leq T$ do
2:     for $i=1,2,\cdots,M$ do
3:        Select action $a^{t}_{i}$ according to $Q^{t}_{i}$ (e.g., $\epsilon$-greedy)
4:        Apply action $a^{t}_{i}$ to the $i$-th arm, observe reward $r^{t}_{i}$ and the new state $s^{t+1}_{i}$
5:        $\delta_{i}^{t}=r^{t}_{i}-\lambda^{t}a^{t}_{i}+\max_{a}Q_{i}^{t}(s^{t+1}_{i},a)-Q_{i}^{t}(s^{t}_{i},a^{t}_{i})-g^{t}_{i}$
6:        $Q_{i}^{t+1}(s^{t}_{i},a^{t}_{i})=Q_{i}^{t}(s^{t}_{i},a^{t}_{i})+\beta^{t}_{i}\delta_{i}^{t}$
7:        $b^{t+1}_{i}=\arg\max_{a}Q_{i}^{t}(s^{t+1}_{i},a)$
8:        $\sigma^{t}_{i}=a^{t}_{i}+D_{i}^{t}(s^{t+1}_{i},b^{t+1}_{i})-D_{i}^{t}(s^{t}_{i},a^{t}_{i})-h^{t}_{i}$
9:        $D_{i}^{t+1}(s^{t}_{i},a^{t}_{i})=D_{i}^{t}(s^{t}_{i},a^{t}_{i})+\alpha^{t}_{i}\sigma^{t}_{i}$
10:     end for
11:     $y^{t}=N-\sum_{i=1}^{M}h_{i}^{t}$
12:     $\lambda^{t+1}=\lambda^{t}-\theta^{t}\mathbf{1}\{|y^{t}|<|y^{t-1}|\}y^{t}$
13:     $t=t+1$
14:  end while
15:  for $i=1,2\cdots,M$ do
16:     for $s\in\mathcal{S}_{i}$ do
17:        $W_{i}(s)=Q_{i}^{T}(s,1)-Q_{i}^{T}(s,0)$ // gain index
18:     end for
19:  end for
Discussion
The definition of gain indices is dependent on $\lambda^{*}$, and our GINO-Q learning aims to learn $\lambda^{*}$ and the associated Q function of every single-arm problem $J_{i}(\lambda^{*})$. This section discusses the robustness of GINO-Q learning to the value of $\lambda^{*}$. Remarkably, the asymptotic optimality of the gain index policy is assured as long as the sequence $\{\lambda^{t}\}$ converges to a neighbourhood of $\lambda^{*}$—not necessarily converging precisely to $\lambda^{*}$.
Formally, for any $\lambda$ and any arm $\mathcal{B}_{i}$, define
$W^{\lambda}_{i}(s)\triangleq Q_{i}(s,1,\lambda)-Q_{i}(s,0,\lambda),\ \forall s\in\mathcal{S}_{i}.$
Then the gain index is $W_{i}(s)=W^{\lambda^{*}}_{i}(s)$. Learning the values of $\{W^{\lambda}_{i}(s)\}$ for a given $\lambda$ constitutes a well-studied reinforcement learning task. However, a practical concern arises: how does the performance of the index policy get affected by inaccuracies in estimating $\lambda^{*}$? Fortunately, the following lemma provides an answer to this question.
Lemma 2.
There exists a non-empty interval $(\lambda^{l},\lambda^{u})$ that contains $\lambda^{*}$, such that the gain index policy with indices $\{W^{\lambda}_{i}(s)\}$ is asymptotically optimal for any $\lambda\in(\lambda^{l},\lambda^{u})$.
The proof of Lemma 2 is provided in the supplementary Appendix. As demonstrated in the proof, $(\lambda^{l},\lambda^{u})$ is the interval where $f^{\prime}(\lambda)=0$, whenever such an interval exists. If no such interval exists, $(\lambda^{l},\lambda^{u})$ corresponds to the range of the two segments connected to $\lambda^{*}$ (as shown in Figure 1). In this latter case, the non-smoothness of $f(\lambda)$ may pose a challenge when searching for $\lambda^{*}$ using gradient decent methods. Specifically, the sequence $\{\lambda^{t}\}$ may oscillate around $\lambda^{*}$ instead of converging to it precisely because $|f^{\prime}(\lambda)|$ is bounded away from 0 for $\lambda$ in the neighborhood of $\lambda^{*}$. Fortunately, according to Lemma 2, this situation will not affect the asymptotic optimality of the resulting index policy. This property makes GINO-Q particularly well-suited for large-scale RMABs—when $M$ is large (with relatively small number of classes), GINO-Q can learn a near-optimal policy and remains robust to the inaccuracies in $\lambda^{*}$.
Figure 1: An illustration of $f(\lambda)$ and $(\lambda^{l},\lambda^{u})$.
Figure 2: Transition probabilities of a nonindexable arm. Except for state 1, each of the remaining states has identical transition probabilities for both actions. For simplicity, we only plot the transitions of action 0 for those states.
Experiments
In this section, we showcase the performance of GINO-Q learning by evaluating it across three distinct RMAB problems. For each problem, we explore various settings characterized by different pairs of $(M,N)$ values, where $M$ is the number of arms, and $N$ denotes the number of arms to be activated at each time.
Baseline algorithms include DQN (Mnih et al. 2015), WIBQ (Avrachenkov and Borkar 2022), and Neurwin (Nakhleh et al. 2021). Among these, DQN treat the entire RMAB as a single MDP. Under this perspective, the state space grows exponentially in $M$, and the number of valid actions is ($M\choose N$. As $M$ increases, the RMAB rapidly becomes an MDP with large discrete state and action spaces, a scenario that DQN struggles to handle. Therefore, we only evaluate DQN in small-scale problems. On the other hand, WIBQ and Neurwin are state-of-the-art algorithms specifically designed for RMABs, with a focus on learning the Whittle index. They serve as our primary baselines for comparison. Additional details of the experiments can be found in the supplementary appendix.
(a) Gain index over training.
(b) Reward over training ($M=10,N=7$).
(c) $M=100,N=70$.
Figure 3: Performances of the GINO-Q and baseline algorithms in a non-indexable RMAB problem. In (a), GI($s$) denotes the gain index of state $s$. States $3$ to $6$ share the same gain index, hence we only plot GI(3) and omit the others. In (b), WIBQ($x$) corresponds to the result of a single run where the Whittle index of state 1 learned by WIBQ is $x\in\{-4,2\}$. In (c), the results represents the average of 20 independent runs, with the shaded areas indicating confidence bounds.
Non-indexable RMAB
Let us first compare the proposed GINO-Q learning with the Whittle-index-based learning algorithms in a non-indexable RMAB. Both WIBQ and Neurwin assume that the given RMAB is indexable because the Whittle index policy is only defined for indexable RMABs. However, determining the indexability of an RMAB can be challenging, particularly in the absence of system knowledge. What happens if we apply a Whittle-index-based learning algorithm to a non-indexable problem? To answer this question, we constructed a non-indexable RMAB and assessed the performances of GINO-Q and WIBQ in this setting. Each arm of this RMAB has 6 states and the transition probabilities among states are as illustrated in Figure 2. The reward function is defined as $r(1,1)=-10,r(1,0)=-4,r(2,1)=r(2,0)=4$ and $r(s,1)=0,r(s,0)=-2$ for $s\in\{3,4,5,6\}$. Let $\mathcal{E}_{\lambda}$ denote the set of states in which action 0 is optimal for the single-arm problem $J_{i}(\lambda)$. It can be verified that $1\in\mathcal{E}_{\lambda}$ for $-4\leq\lambda\leq 2$ and $1\notin\mathcal{E}_{\lambda}$ for $\lambda<-4$ and $\lambda>2$. Hence the set $\mathcal{E}_{\lambda}$ does not expand monotonically as $\lambda$ increases, implying that the arm is not indexable.
We first consider the setting with $M=10$ and $N=7$. Figure 3(a) demonstrates that GINO-Q effectively acquires the gain indices. It is important to note that the key point of the gain index policy is the relative ordering of states by their gain indices. Consequently, the performance of GINO-Q stabilizes once this order is established (compare Figure 3(a) with 3(b)). As discussed above, there are two values of $\lambda$ (i.e., $-4$ and $2$) that satisfy the definition of Whittle index for state 1, while each of the remaining states has a unique Whittle index. In our experiments, WIBQ consistently learned the true Whittle indices for states 2 to 6. However, for state 1, the Whittle index converged to $-4$ in some runs and to $2$ in others. State 1 has the highest priority with index 2 and the lowest priority with index $-4$, resulting in significantly different performances, as shown in Figure 3(b). For benchmarking purposes, we also evaluated the performance of DQN and the random policy (i.e., selecting $N$ arms randomly at each step). Notably, WIBQ performs even worse than the random policy when the Whittle index of state 1 converges to $2$, highlighting the risk of applying Whittle-index-based learning to non-indexable RMABs. For simplicity, the results of Neurwin are not plotted in this setting as WIBQ already learned the true Whittle indices.
We also compared the algorithms in the setting of $(M,N)=(100,70)$, as depicted in Figure 3(c). The low average of WIBQ suggests that it is more likely to converge to $2$ than $-4$. While Neurwin generally outperforms WIBQ in this RMAB, it similarly faces a high risk of learning a very poor index policy. In contrast, GINO-Q consistently learns a near-optimal index policy. Moreover, we computed upper bounds for both settings, provided by the optimal value of the relaxed RMAB. Figure 3(c) shows that when $M$ is large, GINO-Q closely approaches the upper bound, providing strong evidence of its asymptotic optimality.
Figure 4: Performance comparisons between GINO-Q and baseline algorithms in the channel allocation problem.
Figure 5: Performance comparisons between GINO-Q and baseline algorithms in the patrol scheduling problem.
Channel Allocation in Communication Networks
Channel allocation in communication networks involves strategically assigning communication channels to users to optimize the overall system performance. In a simple scenario, a network possesses a finite number of channels, each of which can be allocated to a single user at any time step. A recent research topic is channel allocation aimed at minimizing the average age of information (AoI). AoI is a metric that measures information freshness, defined as the time elapsed since a user last received a message. Tripathi and Modiano (2019) formulated this problem as an RMAB and proved that this RMAB is indexable.
In this experiment, we consider the min-AoI channel allocation problem with two user classes. Each class has a different level of urgency for receiving fresh information, leading us to define a unique cost function of AoI for the class. Let $x$ denote the AoI of a user. The cost functions for the two classes are defined as $f_{1}(x)=2x$ and $f_{2}(x)=\log x$, respectively. The objective is to minimize the overall cost across the network.
We evaluated the GINO-Q and baseline algorithms across three different scales of $M$ and $N$. The reported results represent averages from 20 independent runs, as presented in Figure 4. It can be observed that GINO-Q consistently achieves the best performance across all settings. While Neurwin is capable of learning a Whittle index policy comparable to the gain index policy, it converges at a significantly slower rate. Additionally, compared to Neurwin and WIBQ, GINO-Q exhibits very low variance, as indicated by their confidence bounds.
Patrol Scheduling
Patrol scheduling is another application that is often formulated as an RMAB (Xu et al. 2021). Consider $N$ agents responsible for patrolling $M$ sites. Each agent can visit one site at a time to check for an event of interest. The event is represented as a binary Markov process, where state 1 indicates the event occurs, and state 0 indicates otherwise. The event of interest in a site could be, for example, whether there are illegal intruders in the area. When an agent patrols a site where the event occurs, the controller receives a reward. Conversely, if an event occurs at a site but no agent patrols it, a punishment is generated. In other cases, the controller receives neither reward nor punishment.
Clearly, each site corresponds to an arm in the RMAB. We define two classes of sites, each with a distinct transition matrix for the Markov event. The performances of GINO-Q and the baseline algorithms in this problem are reported in Figure 5. Unfortunately, the upper bounds obtained from the relaxed RMAB appears to be loose for these settings, so they are not plotted here. Nevertheless, once again, GINO-Q outperforms all benchmark algorithms across all settings. In the small-scale setting ($M=6$), DQN performs as well as GINO-Q. However, as $M$ increases to 10, DQN’s policy becomes less effective than both the gain and Whittle index policies due to the curse of dimensionality. When $M=100$, the RMAB exhibits enormous state and action spaces that DQN cannot handle. In contrast, GINO-Q is able to learn a near-optimal index policy quickly, even for large-scale RMABs. Its convergence speed is not affected by the increase in $M$, as long as the number of classes is fixed.
Conclusion
In this paper, we introduced GINO-Q, a novel three-timescale stochastic approximation algorithm designed to address the challenges posed by RMABs. Our approach effectively tackles the curse of dimensionality by decomposing the RMAB into manageable single-arm problems, ensuring that the computational complexity grows linearly with the number of arms. Unlike existing Whittle-index-based learning algorithms, GINO-Q does not require RMABs to be indexable, significantly broadening its applicability. We showed experimentally that Whittle-index-based learning algorithms can perform poorly in non-indexable RMABs. In contrast, GINO-Q consistently learns near-optimal policies across all experimental RMABs, including non-indexable ones, and shows great efficiency by converging significantly faster than existing baselines.
Appendix
Proof of Lemma 2.
Theorem 1 establishes the asymptotic optimality by demonstrating that, as $M\to\infty$, the gain index policy achieves an upper bound on the optimal value of the RMAB. This upper bound corresponds to the optimal value of the relaxed RMAB problem, which is defined by objective (1) subject to constraint (3). We will refer to the optimal policy for the relaxed RMAB problem as the OR policy. Let $x^{t}$ represent the number of arms activated by the OR policy at time $t$. The proof relies on two key facts: (i) when $x^{t}=N$, both the OR policy and the gain index policy activate the same set of arms at time $t$; (ii) as $M\to\infty$, $x^{t}/M$ converges to $N/M$. These facts suggest that in the asymptotic scenario, the gain index policy performs nearly identically to the OR policy. Interested readers can refer to (Chen and Liew 2024) for a detailed proof.
Let $\mu_{\lambda}$ denote the policy for the relaxed RMAB that activates the $i$-th arm according to policy $\pi^{\lambda}_{i}$. Then the OR policy corresponds to policy $\mu_{\lambda^{*}}$. In this proof, we show the existence of a non-empty interval $(\lambda^{l},\lambda^{u})$ such that for any $\lambda$ within this interval: (1) $\mu_{\lambda}$ is identical to $\mu_{\lambda^{*}}$ ; and (2) policy $\mu_{\lambda}$ and the gain index policy defined by $\{W^{\lambda}_{i}(s)\}$ share the relationship described by facts (i) and (ii). This means that in the asymptotic scenario, the gain index policy defined by $\{W^{\lambda}_{i}(s)\}$ performs nearly identically to the OR policy, thereby proving the asymptotic optimality.
Consider the following definition
$\mathcal{E}^{\lambda}_{i}\triangleq\{s\in\mathcal{S}_{i}:W^{\lambda}_{i}(s)>0\},\ i\in[M].$
For the $i$-th arm, $\mathcal{E}^{\lambda}_{i}$ is the set of states in which policy $\pi^{\lambda}_{i}$ takes action 1. Now, for any $\lambda$, if policy $\mu_{\lambda}$ activates exactly $N$ arms at some time $t$, then these $N$ activated arms have positive indices $W^{\lambda}_{i}(s)$, while the remaining arms have non-positive indices at that time. As a result, the index policy with indices $\{W^{\lambda}_{i}(s)\}$ will activate the $N$ arms with positive indices, as they occupy the top $N$ positions. In summary, when policy $\mu_{\lambda}$ activates $N$ arms, the index policy with indices $\{W^{\lambda}_{i}(s)\}$ will match the arm selection of policy $\mu_{\lambda}$.
As stated in Lemma 1, $f(\lambda)$ is a piecewise linear and convex function, and there exists a bounded $\lambda^{*}$ that achieves the minimum value of $f(\lambda)$. Then it is easy to see that $\lambda^{*}$ satisfies one of the following conditions: (1) $f^{\prime}(\lambda^{*})=0$; (2) $f^{\prime}(\lambda^{*})$ is not defined and
$\lim\limits_{\lambda\uparrow\lambda^{*}}f^{\prime}(\lambda)<0,\ \lim\limits_{\lambda\downarrow\lambda^{*}}f^{\prime}(\lambda)>0.$
We start from case (2), $\lambda^{*}$ is the common endpoint of two segments, as shown in fig. 1 of the paper. Let $\lambda^{l}$ denote the left endpoint of the left segment, $\lambda^{u}$ denote the right endpoint of the right segment. Since $g_{i}(\lambda)$ is convex for all $i$, $f^{\prime}(\lambda)=N-\sum_{i=1}^{M}g^{\prime}_{i}(\lambda)$ remains constant within interval $(\lambda^{l},\lambda^{*})$ implies that $\pi^{\lambda}_{i}$ remains unchanged within this interval for all $i$. Using the same argument yields that $\pi^{\lambda}_{i}$ remains unchanged within interval $(\lambda^{*},\lambda^{u})$ for all $i$. Let $\mu^{-}$ and $\mu^{+}$ denote policies $\mu_{\lambda}$ for $\lambda\in(\lambda^{l},\lambda^{*})$ and $\lambda\in(\lambda^{*},\lambda^{u})$, respectively. Then $\mu^{-}$ and $\mu^{+}$ are equally optimal for the relaxed RMAB. Hence $x^{t}/M$ converges to $N/M$ as $M\to\infty$ for both policies.
For the case of $f^{\prime}(\lambda^{*})=0$, there must exists an interval $(\lambda^{l},\lambda^{u})$ such that $f^{\prime}(\lambda)=0$ for $\lambda\in(\lambda^{l},\lambda^{u})$. According to the same argument as before, $\pi^{\lambda}_{i}$ remains unchanged within the interval $(\lambda^{l},\lambda^{u})$ and it is the OR policy.
In summary, there exists a non-empty interval $(\lambda^{l},\lambda^{u})$ containing $\lambda^{*}$ such that, for any $\lambda$ within this interval, the index policy with indices $\{W^{\lambda}_{i}(s)\}$ performs nearly identically to the OR policy in the asymptotic scenario. Hence it is asymptotic optimal. The remaining proof follows standard techniques; for further details, refer to (Chen and Liew 2024). ∎